Logiweb bibliography of prop calc

Up Help

[0] prop calc

[1] base

[2] check

The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-05.UTC:08:46:27.673984 = MJD-53526.TAI:08:46:59.673984 = LGT-4624678019673984e-6